# InvestESG: A multi-agent reinforcement learning benchmark for studying climate investment as a social dilemma

### v1

@@ Abstract
InvestESG is a novel multi-agent reinforcement learning (MARL) benchmark designed to study the impact of Environmental, Social, and Governance (ESG) disclosure mandates on corporate climate investments [^1] Github repo: https://github.com/yuanjiayiy/InvestESG. Supported by both PyTorch and GPU-accelerated JAX framework, the benchmark models an intertemporal social dilemma where companies balance short-term profit losses from climate mitigation efforts and long-term benefits from reducing climate risk, while ESG-conscious investors attempt to influence corporate behavior through their investment decisions. Companies allocate capital across mitigation, greenwashing, and resilience, with varying strategies influencing climate outcomes and investor preferences. Our experiments show that without ESG-conscious investors with sufficient capital, corporate mitigation efforts remain limited under the disclosure mandate. However, when a critical mass of investors prioritizes ESG, corporate cooperation increases, which in turn reduces climate risks and enhances long-term financial stability. Additionally, providing more information about global climate risks encourages companies to invest more in mitigation, even without investor involvement. Our findings align with empirical research using real-world data, highlighting MARL’s potential to inform policy by providing insights into large-scale socio-economic challenges through efficient testing of alternative policy and market designs.
@@ 1 Introduction
Climate change poses a persistent threat to global stability, with droughts, storms, fires, and flooding becoming more intense and frequent (Field et al., 2012), leading to disruption of the natural ecosystem and significant impacts on the global economy. Addressing climate change requires coordinated efforts across multiple sectors, particularly from large corporations, which are reportedly responsible for over 70% of global industrial greenhouse gas emissions (Griffin and Heede, 2017). While adaptation—preparing for the inevitable consequences of climate change—tends to be party-specific and often driven by financial incentives, mitigation—reducing emissions—presents a de facto social dilemma (Leibo et al., 2017), where the benefits of reduced emissions are shared globally yet the costs are borne locally (Olson Jr, 1971; Dahlman, 1979; Buchanan and Stubblebine, 2006). As corporations are inherently self-interested, they are unlikely to reduce emissions voluntarily without external incentives or regulations.
@@
Numerous policies have been proposed to address this challenge. Among these, mandatory Environmental, Social, and Governance (ESG) disclosures have recently been hotly debated in the United States. The Securities and Exchange Commission’s (SEC) proposal, which would require publicly traded companies to disclose climate-related risks, mitigation strategies, and greenhouse gas emissions from their operations, has attracted over 15,000 comments, making it one of the most contentious proposals in the SEC’s history (SEC, 2024a; CNBC, 2024). This has resulted in an indefinite delay in enactment of the policy to allow for further discussion (SEC, 2024b). The US is not alone in facing such a pushback; similar delays are unfolding in the European Union and Korea (Bloomberg News, 2024; Korea Economic Daily, 2023). This highlights the need for thorough research to effectively inform the design and implementation of these policies.
@@
Traditional economics and policy research relies on either empirical analysis—which does not enable testing possible new policies (Doshi et al., 2013; Li and Wu, 2020; Krueger et al., 2021)—or theoretical economics models, which are often limited to scenarios with only two agents (e.g., Friedman et al. 2021), or single-period games (e.g., Pástor et al. 2021). In contrast, Multi-Agent Reinforcement Learning (MARL) enables simulating complex interactions between multiple agents over extended time periods, under diverse hypothesized policy settings. Leveraging MARL to address large-scale socio-economic questions is a growing field (Hertz et al., 2023). Prior work has demonstrated the potential of MARL to design effective taxation schemes that enhance both equality and productivity (Zheng et al., 2021), highlighting its relevance for tackling real-world social challenges.
@@
We propose using multi-agent reinforcement learning (MARL) to explore the impact of the ESG disclosure policy. We introduce InvestESG, an open-source MARL benchmark, to examine how profit-driven corporations balance short-term profits with long-term climate investments and whether ESG-informed investor choices influence corporate behavior. The simulation involves two agent types: companies and investors. Companies allocate funds to mitigation, greenwashing, and resilience, while investors decide whom to invest in based on their preferences for financial returns versus ESG benefits. This creates an intertemporal social dilemma (Hughes et al., 2018), where both agent classes must weigh immediate and local costs against long-term and global gains. Using Schelling diagrams, we demonstrate that in a fully profit-driven environment, a social dilemma arises, but with sufficiently ESG-conscious investors with enough capital, climate mitigation becomes optimal for some or all corporations. However, if companies are able to greenwash to cheaply increase ESG scores without genuine mitigation, the environment once again becomes a social dilemma.
@@
More broadly, we demonstrate how a MARL framework can inform policy debates in the field of climate change. Assessing the effectiveness of a policy is inherently challenging, due to the fact that policy experiments are often prohibitively expensive and impractical to conduct, and even when they are feasible, it can be extremely time-consuming. However, the findings revealed by our environment match existing empirical results, suggesting the simulation has high ecological validity. Given the urgency of addressing climate change, our work provides a new vector for studying this problem, creating a simulated environment where a broad range of regulations can be explored and tested efficiently to provide novel insights into the problem. We present InvestESG as a challenging benchmark for the machine learning community; for researchers that are interested in developing MARL algorithms that can solve social dilemmas, InvestESG represents a social dilemma environment that could have real-world impact. We will provide the code for the benchmark and agent baselines in open-source. We aim to show that MARL algorithmic design can inspire real-world actions that can be leveraged to tackle climate change.
@@ 2 Related Work
Conventional economic methods are limited by either generalizability or tractability. Existing ESG disclosure related research in economics, business, and public policy rely on either empirical data (Doshi et al., 2013; Li and Wu, 2020; Krueger et al., 2021) or simplified theoretical models (Polinsky and Shavell, 2012; Kalkanci and Plambeck, 2020; Cho et al., 2019; Pástor et al., 2021; Friedman et al., 2021). Empirical analyses, while grounded in real-world data, struggle with generalizability and testing counterfactual policies. Theoretical models provide formalized equilibria and explore counterfactuals but are limited by tractability, modeling either multiple agents in single-period games (e.g., (Pástor et al., 2021)) or only two agents over limited time periods (e.g., (Friedman et al., 2021)). By proposing a MARL framework, our method overcomes these limitations by enabling the simulation of complex, multi-agent systems over extended time horizons under diverse policy settings, which allows for emergent behaviors and better captures complex socio-economic dynamics among diverse agents (Hertz et al., 2023).
@@
Current MARL benchmarks and social dilemma environments were not designed to model real-world problems. Various MARL benchmarks have been created to study multi-agent coordination and cooperation; however, they are often limited to simplistic particle simulations (Lowe et al., 2017) or videogames (Agapiou et al., 2023; Carroll et al., 2020; Samvelyan et al., 2019), which have little direct real-world implication. Sequential social dilemmas (SSD) (Leibo et al., 2017) are spatially and temporally extended multi-agent environments in which the payoff to an individual agent for defecting is higher, but if all agents defect the payoff is lower. SSDs go beyond traditional game-theoretic environments like Prisoner’s Dilemma because the complexity of the solving the SSD depends on not only addressing the misalignment between individual and collective rationality, but doing so when the negative consequences of short-sighted actions may take a long time to manifest. Prior research has examined methods to promote cooperation in SSDs by incorporating inequity aversion in agents (Hughes et al., 2018), rewarding agents for influencing others’ actions (Jaques et al., 2018), and enabling agents to provide incentives to others (Yang et al., 2020a). Many of these studies are inspired by factors that drive human cooperation in social dilemmas. However, much of this research has been conducted in environments with no direct real-world implications (e.g. Leibo et al. 2017, 2021). In contrast, our environment directly addresses the problem of climate change, and is benchmarked to real-world climate and economic parameters. We hope to encourage researchers interested in addressing social dilemmas to focus on an environment with potential impact on the problem of climate change.
@@
RL benchmarks can be effective in focusing AI research on climate change issues. Learning to Run a Power Network (L2RPN) is a single-agent RL benchmark focused on improving power grid efficiency (Marot et al., 2021). This work has spawned multiple competitions, and it is currently hosted by the Electric Power Research Institute (EPRI) in conjunction with several other energy companies, government agencies, and universities [^2] https://www.epri.com/l2rpn. L2RPN continues to foster innovative and meaningful collaboration across institutions, showing the potential impact of this type of simplified, simulated RL benchmark. However, L2RPN is focused on single-agent RL, whereas we explore a multi-agent, multi-party social dilemma. The only other multi-agent RL climate benchmark we are aware of was proposed by Zhang et al. (2022), and is focused on studying the dynamics of international climate negotiations, by incorporating an integrated assessment model simulating global climate and economic systems. In contrast, our environment, InvestESG, was designed with a focus on a more targeted policy question regarding ESG disclosure mandates, making our approach more directly applicable to an ongoing and highly debated policy issue.
@@ 3 The InvestESG Environment
InvestESG is a novel, lightweight and GPU-efficient MARL benchmark environment designed to evaluate the impact of ESG disclosure mandates on corporate climate investments. We model the environment as an intertemporal social dilemma where company agents face short-term profit losses from climate mitigation efforts, but these actions reduce long-term climate risks, benefiting all. We provide training scripts based on both PyTorch and JAX implementations, allowing for highly efficient parallel computation across GPUs. This dramatically reduces training time and scales seamlessly to larger agent populations, making InvestESG an ideal platform for rapid experimentation and benchmarking in MARL research. The simulation starts in 2021 and runs through 2120, with each period $t$ corresponding to one year. The environment includes two main components: (1) an evolving climate and economic system, and (2) two types of agents: $M$ company agents $\mathcal{C}_{i}$ for $i\in\{1,\dots,M\}$ and $N$ investor agents $\mathcal{I}_{j}$ for $j\in\{1,\dots,N\}$.
Figure 2: Environment dynamics over the course of a single 100 year episode, for the status quo scenario where all agents are only profit-motivated. Mitigation spending (blue curve) is minimal, leading climate risk (yellow curve) to increase over time. Adverse weather event occurrences are shown as dotted lines; red lines indicate multiple adverse events in a single year.
@@
$P_{t}^{e}=\frac{\mu_{e}t}{1+\lambda_{e}U_{t,m}}+P_{0}^{e},\quad\text{for }e\in\{h,p,d\},$ (1)
[…]
$X_{t}=X_{t,h}+X_{t,p}+X_{t,e},\quad\text{where }X_{t,e}\sim\text{Bernoulli}(P_{t}^{e}),\quad\text{for }e\in\{h,p,d\}.$ (2)
In addition to the evolving climate risks, the environment incorporates a baseline economic growth rate $\gamma$, set to 10% by default, aligned with the historical average annual return of the S&P 500 over the past century (Damodaran, 2024). Company agents’ capital levels $K_{t}^{\mathcal{C}_{i}}$ grow at rate $\gamma$ each year, barring climate events. If an adverse event occurs, company agents lose a portion of their total capital according to their respective climate resilience parameter $L_{t}^{\mathcal{C}_{i}}$.
@@
Company Action Space. Each company agent $\mathcal{C}_{i}$ in period $t$ selects actions from a continuous vector $\mathbf{u}_{t}^{\mathcal{C}_{i}}=(u_{t,m}^{\mathcal{C}_{i}},u_{t,g}^{\mathcal{C}_{i}},u_{t,r}^{\mathcal{C}_{i}})$, where $u_{t,m}^{\mathcal{C}_{i}}$ represents the share of capital allocated to mitigation, $u_{t,g}^{\mathcal{C}_{i}}$ to greenwashing, and $u_{t,r}^{\mathcal{C}_{i}}$ to building climate resilience. The action space for each company is defined as a continuous 3-dimensional unit cube, $\mathcal{U}_{t}^{\mathcal{C}_{i}}=[0,1]^{3}$. If the sum of the three ratios exceeds 1 in any period, the company agent is deemed to be overspending, resulting in a negative capital balance and bankruptcy. Mitigation spending directly reduces the system-wide climate risk as explained above. Greenwashing involves deceptive marketing or accounting tactics that allow companies to appear climate-friendly at a low cost without providing any real benefits to society (de Freitas Netto et al., 2020; Yang et al., 2020b). Resilience spending enhances the company’s climate resilience by lowering its vulnerability $L_{t}^{\mathcal{C}_{i}}$, but it does not reduce emissions and therefore does not mitigate system-wide climate risk. In the default setting, we disable greenwashing and resilience spending to focus on testing companies’ mitigation efforts. These options are later enabled to examine their effects, as will become clear in Section 4.
@@
Investor Action Space. Investor agents first select the companies they want to invest in. Specifically, agent $\mathcal{I}_{j}$ in period $t$ selects an action from a binary vector of length $M$, $\mathbf{a}_{t}^{\mathcal{I}_{j}}=(a_{t,1}^{\mathcal{I}_{j}},\dots,a_{t,M}^{\mathcal{I}_{j}})$, corresponding to the $M$ company agents. Each entry $a_{t,i}^{\mathcal{I}_{j}}=1$ indicates that investor $\mathcal{I}_{j}$ invests in company $\mathcal{C}_{i}$ in period $t$, and $a_{t,i}^{\mathcal{I}_{j}}=0$ otherwise. The investor’s action space at time $t$ is thus $\mathcal{A}_{t}^{\mathcal{I}_{j}}=\{0,1\}^{M}$. Once the choices are made, investors capitals are distributed equally among these chosen companies, as will be detailed later in Equation 4.
Modeling ESG Disclosure. With the ESG disclosure mandate in place, each company agent receives an updated ESG score $Q_{t+1}^{\mathcal{C}_{i}}$ in period $t$, calculated as
$Q_{t+1}^{\mathcal{C}_{i}}=u_{t,m}^{\mathcal{C}_{i}}+\beta u_{t,g}^{\mathcal{C}_{i}},$ (3)
@@
where $\beta>1$ indicates that greenwashing is cheaper than genuine mitigation in terms of building an ESG-friendly image. This mirrors simple ESG ratings provided by some agencies, which typically range from 1 to 100 (S&P Global, 2024) or use letter-based ratings (MSCI Inc., 2024).

### v2

@@ Abstract
InvestESG is a novel multi-agent reinforcement learning (MARL) benchmark designed to study the impact of Environmental, Social, and Governance (ESG) disclosure mandates on corporate climate investments [^1] Github repo: https://github.com/yuanjiayiy/InvestESG. Supported by both PyTorch and JAX implementation, the benchmark models an intertemporal social dilemma where companies balance short-term profit losses from climate mitigation efforts and long-term benefits from reducing climate risk, while ESG-conscious investors attempt to influence corporate behavior through their investment decisions, in a scalable and hardware-accelerated manner. Companies allocate capital across mitigation, greenwashing, and resilience, with varying strategies influencing climate outcomes and investor preferences. Our experiments show that without ESG-conscious investors with sufficient capital, corporate mitigation efforts remain limited under the disclosure mandate. However, when a critical mass of investors prioritizes ESG, corporate cooperation increases, which in turn reduces climate risks and enhances long-term financial stability. Additionally, providing more information about global climate risks encourages companies to invest more in mitigation, even without investor involvement. Our findings align with empirical research using real-world data, highlighting MARL’s potential to inform policy by providing insights into large-scale socio-economic challenges through efficient testing of alternative policy and market designs.
@@ 1 Introduction
Climate change poses a persistent threat to global stability, with droughts, storms, fires, and flooding becoming more intense and frequent (Christopher B Field and Dahe, 2012), leading to disruption of the natural ecosystem and significant impacts on the global economy. Addressing climate change requires coordinated efforts across multiple sectors, particularly from large corporations, which are reportedly responsible for over 70% of global industrial greenhouse gas emissions (Griffin and Heede, 2017). While adaptation—preparing for the inevitable consequences of climate change—tends to be party-specific and often driven by financial incentives, mitigation—reducing emissions—presents a social dilemma (Leibo et al., 2017), where the benefits of reduced emissions are shared globally yet the costs are borne locally (Olson Jr, 1971; Dahlman, 1979; Buchanan and Stubblebine, 2006). As corporations are inherently self-interested, they are unlikely to reduce emissions voluntarily without external incentives or regulations.
@@
Numerous policies have been proposed to address this challenge. Among these, mandatory Environmental, Social, and Governance (ESG) disclosures have recently been hotly debated in the United States. The Securities and Exchange Commission’s (SEC) proposal, which would require publicly traded companies to disclose climate-related risks, mitigation strategies, and greenhouse gas emissions from their operations, has attracted over 15,000 comments, making it one of the most contentious proposals in the SEC’s history (SEC, 2024a; CNBC, 2024). This has resulted in an indefinite delay in enactment of the policy to allow for further discussion (SEC, 2024b). The U.S. is not alone in facing such pushback; similar delays are unfolding in the European Union and Korea (Bloomberg News, 2024; Korea Economic Daily, 2023). This highlights the need for thorough research to effectively inform the design and implementation of these policies.
@@
Traditional economics and policy research relies on either empirical analysis—which does not enable testing possible new policies (Doshi et al., 2013; Li and Wu, 2020; Krueger et al., 2021)—or theoretical economics models, which are often limited to scenarios with only two agents (e.g., (Friedman et al., 2021)), or single-period games (e.g., (Pázstor et al., 2021)). In contrast, Multi-Agent Reinforcement Learning (MARL) enables simulating complex interactions between multiple agents over extended time periods, under diverse hypothesized policy settings. Leveraging MARL to address large-scale socio-economic questions is a growing field (Hertz et al., 2023). Prior work has demonstrated the potential of MARL to design effective taxation schemes that enhance both equality and productivity (Zheng et al., 2021), highlighting its relevance for tackling social challenges.
@@
We propose using multi-agent reinforcement learning (MARL) to explore the impact of the ESG disclosure policy. We introduce InvestESG, an open-source MARL benchmark, to examine how profit-driven corporations balance short-term profits with long-term climate investments and whether ESG-informed investor choices influence corporate behavior. The simulation involves two agent types: companies and investors. Companies allocate funds to mitigation, greenwashing, and resilience, while investors choose investment portfolios based on their preferences for financial returns versus ESG benefits. This creates an intertemporal social dilemma (Hughes et al., 2018), where both agent classes must weigh immediate and local costs against long-term and global gains. Using Schelling diagrams, we demonstrate that in a fully profit-driven environment, a social dilemma arises, but with sufficiently ESG-conscious investors with enough capital, climate mitigation becomes optimal for some or all corporations. However, if companies are able to greenwash to cheaply increase ESG scores without genuine mitigation, the environment once again becomes a social dilemma.
@@
More broadly, we demonstrate the potential of using a MARL framework to inform policy debates in the field of climate change. Importantly, our results are compatible broadly with the existing theoretical and empirical literature that examines this question, which shows its validity in predicting directional behaviors of rational decision makers. Assessing the effectiveness of a policy is inherently challenging, due to the fact that policy experiments are often prohibitively expensive and impractical to conduct, and even when they are feasible, it can be extremely time-consuming. Given the urgency of addressing climate change, our work provides a new vector for studying this problem, creating a simulated environment where a broad range of regulations can be explored and tested efficiently to provide novel insights into the problem.
Our aim is not to claim that our model fully represents real-world complexity. It remains a "first-principles" model; however, it captures aspects that are challenging to integrate into a single framework using traditional economic approaches. We present InvestESG as a challenging benchmark for the machine learning community; for researchers that are interested in developing MARL algorithms that can solve social dilemmas, InvestESG represents a social dilemma environment that could inform policymaking. We will provide the code for the benchmark and agent baselines in open-source.
@@ 2 Related Work
Conventional economic methods are limited by either generalizability or tractability. Existing ESG disclosure related research in economics, business, and public policy rely on either empirical data (Doshi et al., 2013; Li and Wu, 2020; Krueger et al., 2021) or simplified theoretical models (Polinsky and Shavell, 2012; Kalkanci and Plambeck, 2020; Cho et al., 2019; Pázstor et al., 2021; Friedman et al., 2021). Empirical analyses, while grounded in real-world data, struggle with generalizability and testing counterfactual policies. Theoretical models provide formalized equilibria and explore counterfactuals but are limited by tractability, modeling either multiple agents in single-period games (e.g., (Pázstor et al., 2021)) or only two agents over limited time periods (e.g., (Friedman et al., 2021)). By proposing a MARL framework, our method overcomes these limitations by enabling the simulation of complex, multi-agent systems over extended time horizons under diverse policy settings, which allows for emergent behaviors and better captures complex socio-economic dynamics among diverse agents (Hertz et al., 2023).
@@
Current MARL benchmarks and social dilemma environments were not designed to model specific policy problems. Various MARL benchmarks have been created to study multi-agent coordination and cooperation; however, they are often limited to simplistic particle simulations (Lowe et al., 2017) or videogames (Agapiou et al., 2023; Carroll et al., 2020; Samvelyan et al., 2019; Bettini et al., 2024a, b), which have little direct real-world implication. Sequential social dilemmas (SSD) (Leibo et al., 2017) are spatially and temporally extended multi-agent environments in which the payoff to an individual agent for defecting is higher, but if all agents defect the payoff is lower. SSDs go beyond traditional game-theoretic environments like Prisoner’s Dilemma because the complexity of the solving the SSD depends on not only addressing the misalignment between individual and collective rationality, but doing so when the negative consequences of short-sighted actions may take a long time to manifest. Prior research has examined methods to promote cooperation in SSDs by incorporating inequity aversion in agents (Hughes et al., 2018), rewarding agents for influencing others’ actions (Jaques et al., 2018), enabling agents to provide incentives to others (Yang et al., 2020a), and controlling for behavioral diversity (Bettini et al., 2024b). Many of these studies are inspired by factors that drive human cooperation in social dilemmas. However, much of this research has been conducted in environments with no direct real-world implications (e.g. (Leibo et al., 2017, 2021)). In contrast, our environment directly addresses the problem of climate change, and is designed with critical trade-offs around a certain policy-making question. We hope to encourage researchers interested in addressing social dilemmas to focus on an environment with potential impact on the problem of climate change (Bisaro and Hinkel, 2016).
@@
RL benchmarks can be effective in focusing AI research on climate change issues. Learning to Run a Power Network (L2RPN) is a single-agent RL benchmark focused on improving power grid efficiency (Marot et al., 2021). This work has spawned multiple competitions, and it is currently hosted by the Electric Power Research Institute (EPRI) in conjunction with several other energy companies, government agencies, and universities [^2] https://www.epri.com/l2rpn. L2RPN continues to foster innovative and meaningful collaboration across institutions, showing the potential impact of this type of simplified, simulated RL benchmark. However, L2RPN is focused on single-agent RL, whereas we explore a multi-agent, multi-party social dilemma. The only other MARL climate benchmark we are aware of was proposed by Zhang et al. (2022), and is focused on studying the dynamics of international climate negotiations, by incorporating an integrated assessment model simulating global climate and economic systems. In contrast, our environment, InvestESG, was designed with a focus on a more targeted policy question regarding ESG disclosure mandates, making our approach more directly applicable to an ongoing and highly debated policy issue.
3 Problem Setting
Corporations face two key trade-offs in climate action. The first is between short-term costs and long-term benefits, as climate efforts often require upfront investments with returns realized over time. The second is between self-interest and collective benefit, as emission reductions benefit all, but the costs are borne by those who act, which incentivizes free-riding and discourages collective action. These trade-offs motivate us to conceptualize corporate climate actions as an intertemporal social dilemma (Leibo et al., 2017; Bisaro and Hinkel, 2016), which by definition, involves a conflict between immediate self-interest and longer-term collective interests.
We introduce InvestESG, a MARL benchmark environment designed to evaluate the impact of ESG disclosure mandates on corporate climate investments. The environment is grounded in well-established finance and economics literature that explores similar questions, providing a conceptual framework for studying corporate and investor dynamics in response to ESG policies. Pastor et al. (2021) and Pedersen et al. (2021) proposed analytical models of single-period equilibria, where companies either choose or are assigned fixed levels of "greenness," and investors optimize portfolios to balance financial returns with preferences for green assets. We follow a similar setting where companies choose mitigation effort levels every period and receive a single metric ESG score that reflects their greenness. Investors derive utility from the profits generated by their portfolios and, based on their level of "ESG-consciousness", may also receive positive utility by investing in firms with high ESG scores. We extend these frameworks by (1) modeling the evolution of firm-investor strategies over the long run, which poses significant analytical and numerical challenges (Pakes and McGuire, 2001), (2) incorporating climate evolution, where risks worsen over time but can be mitigated by firm efforts, and (3) introducing greenwashing (Lyon and Maxwell, 2011) where companies cheaply inflate ESG scores without genuine mitigation, which raises issues of information asymmetry and principal-agent problems (Grossman and Hart, 1992).
We aim to address the following research questions: (1) Do ESG-conscious investors incentivize companies to undertake mitigation efforts, and how does the level of ESG-consciousness influence these outcomes? (2) Does the presence of heterogeneous investors lead to bifurcation in agent strategies, where only some companies engage in mitigation and attract investments from ESG-conscious investors, while others and their investors opt to free-ride? (3) Are companies likely to engage in greenwashing? (4) What measures can enhance the effectiveness of ESG disclosure mandates?
@@ 4 The InvestESG Environment
In this section, we provide a brief overview of the environment setup. The detailed mathematical formulation is presented in Appendix 9.
The simulation starts in 2021 and runs through 2120, with each period $t$ corresponding to one year. The environment includes two main components: (1) an evolving climate and economic system, and (2) two types of agents: $M$ company agents $\mathcal{C}_{i}$ for $i\in\{1,\dots,M\}$ and $N$ investor agents $\mathcal{I}_{j}$ for $j\in\{1,\dots,N\}$.
(a) Environment dynamics over the course of a single 100 year episode.
(b) Average number of severe climate event occurrences over 10-year period.
Figure 2: Status quo scenario where all agents are only profit-motivated. In (a), mitigation spending (blue curve) is minimal, leading climate risk (yellow curve) to increase over time. Adverse weather event occurrences are shown as dotted lines; red lines indicate multiple adverse events in a single year. In (b), increasing climate risk leads to more frequent occurrence of severe climate events over time.
@@
Company Action Space. Each company agent $\mathcal{C}_{i}$ in period $t$ selects actions from a continuous vector $\mathbf{u}_{t}^{\mathcal{C}_{i}}=(u_{t,m}^{\mathcal{C}_{i}},u_{t,g}^{\mathcal{C}_{i}},u_{t,r}^{\mathcal{C}_{i}})$, where $u_{t,m}^{\mathcal{C}_{i}}$ represents the share of capital allocated to mitigation, $u_{t,g}^{\mathcal{C}_{i}}$ to greenwashing, and $u_{t,r}^{\mathcal{C}_{i}}$ to building climate resilience. The action space for each company is defined as a continuous 3-dimensional unit cube, $\mathcal{U}_{t}^{\mathcal{C}_{i}}=[0,1]^{3}$. If the sum of the three ratios exceeds 1 in any period, the company agent is deemed to be overspending, resulting in bankruptcy. Mitigation spending directly reduces the system-wide climate risk as explained above. Greenwashing involves deceptive marketing or accounting tactics that allow companies to appear climate-friendly at a low cost without providing any real benefits to society (de Freitas Netto et al., 2020; Yang et al., 2020b). Resilience spending enhances the company’s climate resilience by lowering its vulnerability $L_{t}^{\mathcal{C}_{i}}$, but it does not reduce emissions and therefore does not mitigate system-wide climate risk. In the default setting, we disable greenwashing and resilience spending to focus on testing companies’ mitigation efforts. These options are later enabled to examine their effects, as will become clear in Section 5.
@@
Investor Action Space. Investor agents first select the set of companies they want to invest in. Specifically, agent $\mathcal{I}_{j}$ in period $t$ selects an action from a binary vector of length $M$, $\mathbf{a}_{t}^{\mathcal{I}_{j}}=(a_{t,1}^{\mathcal{I}_{j}},\dots,a_{t,M}^{\mathcal{I}_{j}})$, corresponding to the $M$ company agents. Each entry $a_{t,i}^{\mathcal{I}_{j}}=1$ indicates that investor $\mathcal{I}_{j}$ invests in company $\mathcal{C}_{i}$ in period $t$, and $a_{t,i}^{\mathcal{I}_{j}}=0$ otherwise. The investor’s action space at time $t$ is thus $\mathcal{A}_{t}^{\mathcal{I}_{j}}=\{0,1\}^{M}$. Once the choices are made, investors capitals are distributed equally among these chosen companies.
@@
Modeling ESG Disclosure. With the ESG disclosure mandate in place, each company agent receives an updated ESG score $Q_{t+1}^{\mathcal{C}_{i}}$ in period $t$, calculated as $Q_{t+1}^{\mathcal{C}_{i}}=u_{t,m}^{\mathcal{C}_{i}}+\beta u_{t,g}^{\mathcal{C}_{i}}$, where $\beta>1$ indicates that greenwashing is cheaper than genuine mitigation in terms of building an ESG-friendly image. This mirrors simple ESG ratings provided by some agencies, which typically range from 1 to 100 (S&P Global, 2024) or use letter-based ratings (MSCI Inc., 2024).